# STONE: A Submodular Optimization Framework for Active 3D Object Detection

### v1

@@ Abstract
3D object detection is fundamentally important for various emerging applications, including autonomous driving and robotics. A key requirement for training an accurate 3D object detector is the availability of a large amount of LiDAR-based point cloud data. Unfortunately, labeling point cloud data is extremely challenging, as accurate 3D bounding boxes and semantic labels are required for each potential object. This paper proposes a unified active 3D object detection framework, for greatly reducing the labeling cost of training 3D object detector. Our framework is based on a novel formulation of submodular optimization, specifically tailored to the problem of active 3D object detection. In particular, we address two fundamental challenges associated with active 3D object detection: data imbalance and the need to cover the distribution of the data, including LiDAR-based point cloud data of varying difficulty levels. Extensive experiments demonstrate that our method achieves state-of-the-art performance with high computational efficiency compared to existing active learning methods.
@@ 1 Introduction
In many emerging applications such as autonomous driving, it is critical to localize objects in a 3D scene for accurate scene understanding [8, 53]. This is usually achieved by using a 3D detection model based on LiDAR-based point cloud data with oriented bounding boxes and semantic labels. While highly accurate recognition and localization of objects can be achieved due to recent advancements in deep learning, this performance often comes at the expense of requiring a large volume of labeled point cloud data, which is much more costly to collect compared to typical RGB images [48, 49].
Active learning (AL) is the standard method for reducing labeling costs in machine learning [7, 42]. AL often starts with a small labeled set and iteratively selects the most informative samples for label acquisition from a large pool of unlabeled data, given a labeling budget. The informativeness, of a selected sample, can be measured in various ways. For example, by computing the sampling uncertainty as measured by maximum Shannon entropy [45, 31] or estimated model changes or finding the most representative samples to avoid sample redundancy [35, 17, 38] by using greedy coreset algorithms [41, 18] or clustering-based approaches [37, 54, 3].
@@
In a recent study, CRB [34] proposed three stages to label unlabeled point clouds hierarchically, ensuring they are concise, representative, and geometrically balanced. One of the critical components of CRB [34] is to achieve label balance in individual point clouds. However, it cannot guarantee the label distribution is balanced across all the labeled point clouds. More recently, KECOR [33] was proposed for characterizing sample informativeness in both classification and regression tasks with a unified measurement. The main idea is to select a subset of point clouds that maximizes the kernel coding rate. This approach enables the selection of representative samples from the unlabeled point cloud. However, it does not address the issue of data imbalance.
In this paper, we propose a unified submodular optimization framework, called STONE, for active 3D object detection, to address limitations in existing methods. Our framework leverages submodular functions [12], a classical tool for measuring set quality, due to its well-known property of diminishing returns. Building on these fundamental results, we introduce a novel formulation of submodular optimization for active 3D object detection. This formulation not only ensures that the selected point cloud achieves maximal coverage of the unlabeled point cloud but also addresses the issue of data imbalance. Based on the formulation, we propose a two-stage algorithm. Firstly, we select representative point clouds using a submodular function based on gradients computed from hypothetical labels using Monte Carlo dropout [13]. Secondly, we employ a greedy search algorithm to select unlabeled point clouds, aiming to balance the data distribution as measured by another submodular function. Our work makes the following contributions. 1) We introduce the first submodular optimization framework for active 3D object detection, addressing two fundamental challenges in this domain. 2) We develop a simple and efficient two-stage algorithm within the framework for selecting representative point clouds and addressing the issue of data imbalance. 3) We extensively validate the proposed framework on real-world autonomous driving datasets, including KITTI [15] and Waymo Open [49], achieving state-of-the-art performance in active 3D object detection. Furthermore, additional results in active 2D object detection demonstrate the high generalizability of our proposed method.
@@ 2 Related Works
Active Learning (AL) has been a deeply studied topic in machine learning [7, 42], which involves alleviating labeling annotation costs by selecting the most representative samples from a pool of unlabelled data, all the while not comprising model performance. Broadly speaking, AL can be categorized into two strategies - uncertainty sampling and representative/diversity sampling. Algorithms under representative/diversity sampling select subsets of data that act as stand-ins or surrogates for the entire dataset [1]. Prior works have explored coreset-based subset selection [41, 18], clustering algorithms [37, 54, 3], and generative adversarial learning [16]. In uncertainty sampling [30], samples are selected by maximizing the Shannon entropy [31] of the posterior probability, minimizing the model’s subsequent training error through variance reduction [6], based on the maximum gradient magnitude [43], and considering the disagreement among a committee of model hypotheses [44, 51]. Hybrid sampling strategies are proposed recently [24, 1, 25, 20], combining representative/diversity and uncertainty sampling, like BADGE [1], where the authors propose to use the gradient magnitude with respect to the classifier’s parameters as a measure of uncertainty. This approach selects samples whose gradients cover a wide range of directions.
Active Learning for 3D Object Detection. While AL has been actively studied and applied to image classification and regression tasks, it has recently been garnering interest in the 3D object detection community. Initial works like MC-MI [11] make use of Monte-Carlo(MC)-dropout [14] and Deep Ensembles [29] to compute the Shannon entropy in the predicted labels and mutual information between class predictions and model parameters. CONSENSUS [40] estimates uncertainty using ensembles for 2D/3D object detection. While these works relied on generic metrics to measure prediction uncertainty, two other works on 3D object detection, CRB [34] and KECOR [33], were proposed recently. In CRB [34], the method greedily searches and samples unlabelled point clouds that exhibit concise labels, representative features, and geometric balance. KECOR [33] shows that, for sample selection, maximizing the kernel coding rate is beneficial and improves over task-specific AL methods for 3D detection, at fast-running times.
@@ 3 Background
Active Learning for 3D Object Detection. Active 3D object detection aims to reduce the labeling cost for training the 3D detector. In the beginning stage, a small number of point clouds $D_{L}$ are randomly selected from the unlabeled data pool $D_{U}$ to train the backbone 3D object detection model. During active learning, for each query iteration $q\in\{1,2,\ldots,Q\}$, a given active learning method will select $\Gamma$ number of unlabeled point clouds from $D_{U}$. These selected point clouds are then given to a human annotator for labeling the bounded boxes and semantic labels, and the labeled point clouds $D_{S}=\{P_{j},B_{j}\}_{j\in[N_{q}]}$ along with the previously selected point clouds are combined to create selected point clouds $D_{L}=D_{L}\cup D_{S}$ for the backbone model to re-train. When the query round $Q$ is reached or the total queried bounding box number, or budget, $N_{Q}$ is reached, it will stop repeating the above process, where $N_{Q}$ = $\sum_{q=1}^{Q}N_{q}$.
@@ 4.1 A Submodular Optimization Approach to Active 3D Object Detection
To achieve the first goal, we leverage a submodular function $f_{1}$ to measure the representativeness of the selected unlabeled point clouds $D_{S}$ with respect to the whole unlabeled set $D_{U}$. This can be achieved by minimizing the absolute difference between $f_{1}(D_{U})$ and $f_{1}(D_{S})$. To achieve the second goal, we leverage a different submodular function $f_{2}$ to ensure that once the selected unlabeled point clouds $D_{S}$ are added to the labeled set $D_{L}$, the overall quality, as measured by label distribution, will not decrease. In sum, we aim to select unlabeled point clouds $D_{S}$ from the unlabeled pool $D_{U}$ that optimize the following two objectives,
@@ 4.2 STONE
In order to optimize the submodular optimization framework in Equation 3, we have designed an active learning pipeline with two stages, aligning with the dual objectives of the submodular criteria. To enhance computational efficiency and simplify optimization, we have implemented a hierarchical structure to eliminate unselected samples at each stage, with the remaining samples comprising our final selection. Initially, we select $\Gamma_{1}$ samples using the proposed Gradient-Based Submodular Subset Selection (GBSSS), which maximizes diversity and coverage while minimizing redundancy from $D_{U}$. Subsequently, we select $\Gamma_{2}$ samples from $\Gamma_{1}$ through Submodular Optimization for Class Balancing (SDMCB). A detailed explanation of these stages is provided in the following paragraphs.
@@ 4.2.1 Gradient-Based Submodular Subset Selection (GBSSS)
However, when the dataset is highly imbalanced, as is usually the case in 3D object detection, the gradients generated by the above approach become inaccurate for classes with fewer samples. To address this issue, we introduce two novel reweighing approaches for handling regression loss $L_{reg}$ and classification loss $L_{cls}$ in 3D object detection for a better computation of the gradients for rare classes. In particular, after calculating the regression loss for each bounding box, we calculate the average loss $L_{reg}^{c}$ for class $c$ based on the regression hypothetical labels. For a given label $c$, the regression loss reweighing factors can be expressed as $w_{c}=\frac{1}{n_{c}}$, where $n_{c}$ is the number of the bounding boxes that belong to class $c$. We then normalize $w_{c}$ as $\tilde{w}_{c}=\frac{w_{c}}{\max(w_{c})}$. The purpose of doing so is two-fold - 1) The class-specific weights in $L_{reg}^{c}$ are uniformly scaled to prevent bias toward frequently occurring classes. 2) This approach helps reduce overfitting to bounding boxes associated with those classes. Then, we perform an element-wise multiplication between $\tilde{w}_{c}$ and $L_{reg}^{c}$ to re-scale the loss, for each class $c$, and compute the mean, across all the semantic classes $C$, to get the final reweighed regression loss $\hat{L}_{reg}$ as $\hat{L}_{reg}=\frac{1}{C}\sum_{c=1}^{C}\tilde{w}_{c}\cdot L_{reg}^{c}$. The reweighed regression loss $\hat{L}_{reg}$ places more emphasis on classes with fewer samples, which is critical for computing the gradient.
@@
where $g(x)=\log(1+x)$. The concavity of $g(x)$ ensures that the marginal gain of adding more instances of similar features decreases as more such instances are added. This means that the function will favor adding elements that introduce new features rather than redundant ones. For the score function [56], $\mu(\cdot)$, we use gradients distribution entropy $H(\nabla_{\boldsymbol{\theta}}\mathcal{L})$ as the measurement. More details are given in the Supplementary. A greedy algorithm is then applied to iteratively add samples to a subset, providing the most significant increase in marginal gain until the target size, $\Gamma_{1}$, is reached.
@@ 4.2.2 Submodular Diversity Maximization for Class Balancing (SDMCB)
One of the main challenges for active 3D object detection is that each 3D scene can contain multiple semantic classes, such as cars and cyclists. As the queried point clouds increase, it will inevitably introduce an imbalanced label distribution in $D_{L}$, eventually resulting in performance drops. This situation is more prominent in autonomous driving datasets. A recent study, CRB [34], attempts to address this imbalance issue by introducing individual point cloud balancing. However, it cannot guarantee a balanced label distribution in $D_{U}$ across multiple active learning queries. In Figure 1, we illustrate the imbalance in label distribution entropy caused by existing methods [34, 33], particularly KECOR [33], where label imbalance begins to increase as the number of query rounds rises. Aiming to alleviate this label imbalance, we utilize entropy calculated from the predicted label distribution as the submodular function $f_{2}$. This approach supports the principle of diminishing returns as $D_{L}$ becomes more balanced. Our method is divided into two steps. In Step 1, we select point clouds which have balanced label distribution based on the hypothetical labels. In Step 2, from the selected ones, we further choose a subset that can be labeled to ensure that the label distribution is balanced in the labeled set $D_{L}$.
@@ 5.1 Experimental Setup
Datasets. For our experiments, we use the KITTI [15] dataset, one of the commonly used datasets in autonomous driving tasks. The dataset consists of 3,712 training samples and 3,769 validation samples, that includes a total of 80,256 labeled objects. These objects include cars, pedestrians, and cyclists, each annotated with class categories and bounding boxes. We also use the more challenging dataset for 3D object detection in autonomous driving - the Waymo Open dataset [49]. It includes 158,361 training samples and 40,077 testing samples. In the KITTI dataset, task difficulty levels are defined as EASY for fully visible objects, MODERATE for partially occluded objects, and HARD for significantly occluded objects. For the Waymo test set, the framework categorizes difficulty into two levels: LEVEL 1 with more than five LiDAR points inside the ground-truth bounding box and LEVEL 2 with less or equal to five points. The sampling intervals are set to 1 and 10 for KITTI [15] and Waymo Open [49], respectively.
@@
Evaluation Metrics. To maintain fairness with the baselines on KITTI [15], we measure the performance using Average Precision (AP) for 3D and Bird Eye View (BEV) detection, with rotated Intersection over Union (IoU) thresholds of 0.7 for cars and 0.5 for pedestrians and cyclists, following [46]. For the Waymo Open dataset [49], performance is measured using Average Precision (AP) and Average Precision Weighted by Heading (APH), with IoU thresholds of 0.7 for vehicles and 0.5 for pedestrians and cyclists.
Implementation Details. We train our proposed method and all baselines on a NVIDIA RTX A5000 GPU. To make fair comparisons, we adopt the implementation settings as outlined in CRB [34]. All the baselines use PV-RCNN [46] as the backbone detection model.
  • Training settings. For KITTI [15] and Waymo Open [49], the training batch sizes are set to 6 and 4 respectively. However, the evaluation batch sizes are set to 16 for both datasets. We optimize the network parameters using Adam with a fixed learning rate of 0.01. For all the methods, we perform 5 stochastic forward passes of the MC-Dropout [14].
  • Active learning parameters. For KITTI [15], we set $\Gamma_{1}$ and $\Gamma_{2}$ to 400 and 300 respectively. In the case of Waymo Open [49], $\Gamma_{1}$ and $\Gamma_{2}$ are set to 2,000 and 1,200 respectively. To ensure fairness, $N_{q}$ is set to 100 in all the methods.

### v2

@@ Abstract
3D object detection is fundamentally important for various emerging applications, including autonomous driving and robotics. A key requirement for training an accurate 3D object detector is the availability of a large amount of LiDAR-based point cloud data. Unfortunately, labeling point cloud data is extremely challenging, as accurate 3D bounding boxes and semantic labels are required for each potential object. This paper proposes a unified active 3D object detection framework, for greatly reducing the labeling cost of training 3D object detectors. Our framework is based on a novel formulation of submodular optimization, specifically tailored to the problem of active 3D object detection. In particular, we address two fundamental challenges associated with active 3D object detection: data imbalance and the need to cover the distribution of the data, including LiDAR-based point cloud data of varying difficulty levels. Extensive experiments demonstrate that our method achieves state-of-the-art performance with high computational efficiency compared to existing active learning methods. The code is available at https://github.com/RuiyuM/STONE
@@ 1 Introduction
In many emerging applications such as autonomous driving, it is critical to localize objects in a 3D scene for accurate scene understanding [8, 56]. This is usually achieved by using a 3D detection model based on LiDAR point cloud data with oriented bounding boxes and semantic labels. While highly accurate recognition and localization of objects can be achieved due to recent advancements in deep learning, this performance often comes at the expense of requiring a large volume of labeled point cloud data, which is much more costly to collect compared to typical RGB images [51, 52].
Active learning (AL) is the standard method for reducing labeling costs in machine learning [7, 45]. AL often starts with a small labeled set and iteratively selects the most informative samples for label acquisition from a large pool of unlabeled data, given a labeling budget. The informativeness, of a selected sample, can be measured in various ways. For example, by computing the sampling uncertainty as measured by maximum Shannon entropy [48, 32] or estimated model changes or finding the most representative samples to avoid sample redundancy [36, 17, 40] by using greedy coreset algorithms [44, 18] or clustering-based approaches [38, 57, 3, 42].
@@
In a recent study, CRB [35] proposed three stages to label unlabeled point clouds hierarchically, ensuring they are concise, representative, and geometrically balanced. One of the critical components of CRB is to achieve label balance in individual point clouds. However, it cannot guarantee the label distribution is balanced across all the labeled point clouds. More recently, KECOR [34] was proposed for characterizing sample informativeness in both classification and regression tasks with a unified measurement. The main idea is to select a subset of point clouds that maximizes the kernel coding rate. This approach enables the selection of representative samples from the unlabeled point cloud. However, it does not address the issue of data imbalance.
In this paper, we propose a unified submodular optimization framework, called STONE, for active 3D object detection, to address limitations in existing methods. Our framework leverages submodular functions [12], a classical tool for measuring set quality, due to its well-known property of diminishing returns. Building on these fundamental results, we introduce a novel formulation of submodular optimization for active 3D object detection. This formulation not only ensures that the selected point cloud achieves maximal coverage of the unlabeled point cloud but also addresses the issue of data imbalance. Based on the formulation, we propose a two-stage algorithm. Firstly, we select representative point clouds using a submodular function based on gradients computed from hypothetical labels using Monte Carlo dropout [13]. Secondly, we employ a greedy search algorithm to select unlabeled point clouds, aiming to balance the data distribution as measured by another submodular function. Our work makes the following core contributions. 1) We introduce the first submodular optimization framework for active 3D object detection, addressing two fundamental challenges in this domain. 2) We develop a simple and efficient two-stage algorithm within the framework for selecting representative point clouds and addressing the issue of data imbalance. 3) We extensively validate the proposed framework on real-world autonomous driving datasets, including KITTI [15] and Waymo Open dataset [52], achieving state-of-the-art performance in active 3D object detection. Furthermore, additional results in active 2D object detection demonstrate the high generalizability of our proposed method.
@@ 2 Related Works
Active Learning (AL) has been a deeply studied topic in machine learning [7, 45], which involves alleviating labeling annotation costs by selecting the most representative samples from a pool of unlabeled data, all the while not comprising model performance. Broadly speaking, AL can be categorized into two strategies - uncertainty sampling and representative/diversity sampling. Algorithms under representative/diversity sampling select subsets of data that act as stand-ins or surrogates for the entire dataset [1]. Prior works have explored coreset-based subset selection [44, 18], clustering algorithms [38, 57, 3], and generative adversarial learning [16]. In uncertainty sampling [31], samples are selected by maximizing the Shannon entropy [32] of the posterior probability, minimizing the model’s subsequent training error through variance reduction [6], based on the maximum gradient magnitude [46], and considering the disagreement among a committee of model hypotheses [47, 54]. Hybrid sampling strategies are proposed recently [25, 1, 26, 20], combining representative/diversity and uncertainty sampling, like BADGE [1], where the authors propose to use the gradient magnitude with respect to the classifier’s parameters as a measure of uncertainty. This approach selects samples whose gradients cover a wide range of directions.
Active Learning for 3D Object Detection. While AL has been actively studied and applied to image classification and regression tasks, it has recently been garnering interest in the 3D object detection community. Initial works like MC-MI [11] make use of Monte-Carlo (MC)-dropout [14] and Deep Ensembles [30] to compute the Shannon entropy in the predicted labels and mutual information between class predictions and model parameters. CONSENSUS [43] estimates uncertainty using ensembles for 2D/3D object detection. While these works relied on generic metrics to measure prediction uncertainty, two other works on 3D object detection, CRB [35] and KECOR [34], were proposed recently. In CRB, the method greedily searches and samples unlabeled point clouds that exhibit concise labels, representative features, and geometric balance. KECOR shows that, for sample selection, maximizing the kernel coding rate is beneficial and improves over task-specific AL methods for 3D detection, at fast-running times.
@@ 3 Background
Active Learning for 3D Object Detection. Active 3D object detection aims to reduce the labeling cost for training the 3D detector. In the beginning stage, a small number of labeled point clouds $D_{L}$ are randomly selected from the unlabeled data pool $D_{U}$ to train the backbone 3D object detection model. During active learning, for each query iteration $q\in\{1,2,\ldots,Q\}$, a given active learning method will select $\Gamma$ number of unlabeled point clouds from $D_{U}$. These selected point clouds are then given to a human annotator for labeling the bounded boxes and semantic labels, and the labeled point clouds $D_{S}=\{P_{j},B_{j}\}_{j\in[N_{q}]}$ along with the previously selected point clouds are combined to create selected point clouds $D_{L}=D_{L}\cup D_{S}$ for the backbone model to re-train. When the query round $Q$ is reached or the total queried bounding box number, or budget, $N_{Q}$ is reached, it will stop repeating the above process, where $N_{Q}$ = $\sum_{q=1}^{Q}N_{q}$.
@@ 4.1 A Submodular Optimization Approach to Active 3D Object Detection
[…]
@@
Figure 1: STONE: An illustrative pipeline of our proposed active learning method for 3D object detection leveraging submodular functions.
@@ 4.2 STONE
In order to optimize the submodular optimization framework in Equation 3, we have designed an active learning pipeline with two stages, aligning with the dual objectives of the submodular criteria. To enhance computational efficiency and simplify optimization, we have implemented a hierarchical structure to eliminate unselected samples at each stage, with the remaining samples comprising our final selection. Initially, we select $\Gamma_{1}$ samples using the proposed Gradient-Based Submodular Subset Selection (GBSSS), which maximizes diversity and coverage while minimizing redundancy from $D_{U}$. Subsequently, we select $\Gamma_{2}$ samples from $\Gamma_{1}$ through Submodular Optimization for Class Balancing (SDMCB). A detailed explanation of these stages is provided in the following paragraphs. We illustrate our proposed pipeline in Figure 1.
@@ 4.2.1 Gradient-Based Submodular Subset Selection (GBSSS)
However, when the dataset is highly imbalanced, as is usually the case in 3D object detection, the gradients generated by the above approach become inaccurate for classes with fewer samples [39]. To address this issue, we introduce two novel reweighing approaches for handling regression loss $L_{reg}$ and classification loss $L_{cls}$ in 3D object detection for a better computation of the gradients for rare classes. In particular, after calculating the regression loss for each bounding box, we calculate the average loss $L_{reg}^{c}$ for class $c$ based on the regression hypothetical labels. For a given label $c$, the regression loss reweighing factors can be expressed as $w_{c}=\frac{1}{n_{c}}$, where $n_{c}$ is the number of the bounding boxes that belong to class $c$. We then normalize $w_{c}$ as $\tilde{w}_{c}=\frac{w_{c}}{\max(w_{c})}$. The purpose of doing so is two-fold - 1) The class-specific weights in $L_{reg}^{c}$ are uniformly scaled to prevent bias toward frequently occurring classes. 2) This approach helps reduce overfitting to bounding boxes associated with those classes. Then, we perform an element-wise multiplication between $\tilde{w}_{c}$ and $L_{reg}^{c}$ to re-scale the loss, for each class $c$, and compute the mean, across all the semantic classes $C$, to get the final reweighed regression loss $\hat{L}_{reg}$, as $\hat{L}_{reg}=\frac{1}{C}\sum_{c=1}^{C}\tilde{w}_{c}\cdot L_{reg}^{c}$. The reweighed regression loss $\hat{L}_{reg}$ places more emphasis on classes with fewer samples, which is critical for computing the gradient.
@@
where $g(x)=\log(1+x)$. The concavity of $g(x)$ ensures that the marginal gain of adding more instances of similar features decreases as more such instances are added. This means that the function will favor adding elements that introduce new features rather than redundant ones. For the score function [59], $\mu(\cdot)$, we use gradients distribution entropy $H(\nabla_{\boldsymbol{\theta}}\mathcal{L})$ as the measure of informativeness. More details are given in the Appendix. A greedy algorithm is then applied to iteratively add samples to a subset, providing the most significant increase in marginal gain until the target size, $\Gamma_{1}$, is reached.
@@ 4.2.2 Submodular Diversity Maximization for Class Balancing (SDMCB)
One of the main challenges for active 3D object detection is that each 3D scene can contain multiple semantic classes, such as cars and cyclists. As the queried point clouds increase, it will inevitably introduce an imbalanced label distribution in $D_{L}$, eventually resulting in performance drops. This situation is more prominent in autonomous driving datasets. A recent study, CRB, attempts to address this imbalance issue by introducing individual point cloud balancing. However, it cannot guarantee a balanced label distribution in $D_{U}$ across multiple active learning queries. In Figure 2, we illustrate the imbalance in label distribution entropy caused by existing methods [35, 34], particularly KECOR, where label imbalance begins to increase as the number of query rounds rises. Aiming to alleviate this label imbalance, we utilize entropy calculated from the predicted label distribution as the submodular function $f_{2}$. This approach supports the principle of diminishing returns as $D_{L}$ becomes more balanced. Our method is divided into two steps. In Step 1, we select point clouds which have balanced label distribution based on the hypothetical labels. In Step 2, from the selected ones, we further choose a subset that can be labeled to ensure that the label distribution is balanced in the labeled set $D_{L}$.
@@ 5.1 Experimental Setup
Datasets. For our experiments, we use the KITTI dataset [15], one of the commonly used datasets in autonomous driving tasks. The dataset consists of 3,712 training samples and 3,769 validation samples, which include a total of 80,256 labeled objects. These objects include cars, pedestrians, and cyclists, each annotated with class categories and bounding boxes. We also use the more challenging dataset for 3D object detection in autonomous driving - the Waymo Open dataset [52]. It includes 158,361 training samples and 40,077 testing samples. In the KITTI dataset, task difficulty levels are defined as EASY for fully visible objects, MODERATE for partially occluded objects, and HARD for significantly occluded objects. For the Waymo test set, the framework categorizes difficulty into two levels: LEVEL 1 with more than five LiDAR points inside the ground-truth bounding box and LEVEL 2 with less or equal to five points. The sampling intervals are set to 1 and 10 for KITTI and Waymo Open, respectively.
@@
Evaluation Metrics. To maintain fairness with the baselines on KITTI, we measure the performance using Average Precision (AP) for 3D and Bird Eye View (BEV) detection, with rotated Intersection over Union (IoU) thresholds of 0.7 for cars and 0.5 for pedestrians and cyclists, following [49]. For the Waymo Open, performance is measured using Average Precision (AP) and Average Precision Weighted by Heading (APH), with IoU thresholds of 0.7 for vehicles and 0.5 for pedestrians and cyclists.
Implementation Details. We train our proposed method and all baselines on a GPU cluster with 4 NVIDIA RTX A5000 GPUs. To make fair comparisons, we adopt the implementation settings as outlined in CRB. All the baselines use PV-RCNN [49] as the backbone detection model.
  • Training settings. For KITTI and Waymo Open, the training batch sizes are set to 6 and 4 respectively. However, the evaluation batch sizes are set to 16 for both datasets. We optimize the network parameters using Adam with a fixed learning rate of 0.01. For all the methods, we perform 5 stochastic forward passes of the MC-Dropout [14].
  • Active learning parameters. For KITTI, we set $\Gamma_{1}$ and $\Gamma_{2}$ to 400 and 300 respectively. In the case of Waymo Open, $\Gamma_{1}$ and $\Gamma_{2}$ are set to 2,000 and 1,200 respectively. To ensure fairness, $N_{q}$ is set to 100 in all the methods.